OpenStudy (brightemmerald):

What number must you add to complete the square? x^2 + 8x = -3 A. 16 B. 8 C. 4 D. 64

OpenStudy (3mar):

May I help?

OpenStudy (brightemmerald):

Of course!

OpenStudy (triciaal):

@3mar are you going to help or not?

OpenStudy (3mar):

I am !

OpenStudy (3mar):

As the coefficient of x^2 is unity, then we are going to add and subtract the half of the x-coefficient squared!

OpenStudy (brightemmerald):

uhh...

OpenStudy (3mar):

That means look at : \[(1)x^2\] then: \[ ( )x\] \[x^2+8x+(\frac{ 8 }{ 2 })^2-(\frac{ 8 }{ 2 })^2=-3\]

OpenStudy (brightemmerald):

I'm still not comprehending

OpenStudy (3mar):

to get \[x^2+8x+16=(x+4)^2\] you must add 16 and subtract 16 easy now?

OpenStudy (brightemmerald):

Partially...

OpenStudy (3mar):

add 16 and subtract 16 don't change anything in one side of the equation!

OpenStudy (brightemmerald):